OpenStudy (anonymous):

What rule represents the function. (0,0)(1,0.5)(2,2)(3,4.5)(4,8)

OpenStudy (anonymous):

look at the average rate of changes: \[ (0.5, 0.5)\\ (1.5,1.5)\\ (2.5,2.5)\\ (3.5,2.5) \]

OpenStudy (anonymous):

That's a sign that is probably an exponential function.

OpenStudy (anonymous):

I would try: \[ f(x)= ab^x+c \]Since \(f(0)=0\):\[ 0=ab^0+c\implies a=-c \]

OpenStudy (anonymous):

So\[ f(x)=ab^x-a \]

OpenStudy (anonymous):

You put \(f(1)=\frac 12\)\[ \frac 12=ab^1-a\implies b=\frac 1{2a}+1 \]

OpenStudy (anonymous):

Keep doing this and you'll get it.

OpenStudy (anonymous):

\[ 2=ab^2-a \]

OpenStudy (anonymous):

@meggersee Do you get it?